### v1

# Jaynes-Cummings model in a unitary fractional-time description

Abstract
The time-evolution operator derived from the fractional-time Schrödinger equation is considered non-unitary because it fails to preserve the norm of the vector state as time evolves. However, considering the time-dependent non-Hermitian quantum formalism to the time-fractional dynamics, it has been demonstrated that a unitary evolution can be achieved for a traceless two-level Hamiltonian. This is accomplished by considering a dynamical Hilbert space embedding a time-dependent metric operator, with respect to which the system evolves in a unitary manner, allowing for the proper interpretation of standard quantum mechanical probabilities. In this work, we apply the unitary description to the Jaynes-Cummings model in the fractional-time scenario for investigating the modification in terms of the fractional-order parameter $\alpha$ of the well-known dynamical properties, such as the atomic population inversion of the two-level atom, and the atom-field entanglement when the atom starts in its excited state and field is initially in a coherent state.
I Introduction
Many developments at the interface of physics and fractional calculus have been drawing attention. The cornerstone of this approach lies in substituting the $n$-order derivative $\partial_{z}^{n}f(z)$ by fractional-order differential operators denoted as $\mathcal{D}^{\alpha}_{z}f(z)$, which represents fractional derivatives of order $\alpha$ acting on the function $f(z)$ with respect to the variable $z$. The specific definition of the operator depends on the underlying mathematical functions involved [1]. Interesting applications of fractional calculus appears in statistical physics, particularly in the context of continuous-time random walks (CTRW) to model transport phenomena. CTRW offer a versatile framework for describing anomalous diffusion processes, commonly observed in complex systems exhibiting memory effects and non-local interactions. In the spatial domain, fractional derivative leads to the emergence of Lévy flights [2]. These are characterized by jumps with power-law distributed step lengths, resulting in a diffusion process with long-tailed probability densities. Conversely, incorporating a fractional derivative in the temporal domain leads to subdiffusive behavior [3]. This signifies slower-than-expected diffusion, often observed in systems with heterogeneous environments or trapping mechanisms. The choice of which domain (spatial or temporal) to introduce the fractional derivative depends on the specific physical mechanisms governing the transport process.
In quantum mechanics, a branch of quantum physics rooted in fractional calculus, has emerged as a powerful framework for understanding the behavior of quantum systems with non-local, non-Markovian, and long-range interactions. This burgeoning field encompasses diverse areas such as Lévy flights over quantum paths [4], optics [5, 6], $\mathcal{PT}$-symmetric systems [7], the nonlinear variable-order time fractional Schrödinger equation [8], disorder in the vibrational spectra [9], time-dependent quantum potentials [10], and anomalous diffusion in three-level system [11]. An experimental demonstration by Wu et al. [12] investigated spontaneous emission from a two-level atom in anisotropic one-band photonic crystals. They elegantly employed fractional calculus to resolve an unphysical bound state anomaly arising when the resonant atomic frequency deviates from the photonic band gap. This anomaly, characterized by an infinitely long lifetime, vanishes when the emission peak aligns with the band gap [13].
In this realm, the means of characterizing states involves the application of the fractional Schrödinger equation (FSE), which was first introduced by Laskin. Unlike the conventional Schrödinger equation, Laskin’s formulation replaces the standard second-order spatial derivative with a fractional Laplacian operator derived from the Reisz derivative [14, 15, 16, 17]. This adaptation allows the FSE to model non-local interactions and memory effects, which are critical for comprehending the nuanced transport processes within quantum systems. Furthermore, Naber [18] has proposed a fractional-time Schrödinger equation (FTSE) assuming the Caputo fractional derivative in the place of the ordinary time derivative in such a way that the equation is written as
$i^{\alpha}\hbar_{\alpha}{}^{~{}\text{C}}_{~{}0}\mathcal{D}_{t}^{\alpha}|\Psi_{\alpha}(t)\rangle=\hat{H}_{\alpha}|\Psi_{\alpha}(t)\rangle\text{,}$ (1)
with
[…]
defining the fractional Caputo derivative for $\alpha\in[0,1)$. $\hat{H}_{\alpha}$ represents the fractional Hamiltonian, and $\hbar_{\alpha}$ the fractional-Planck constant used as a scale factor [see [18]], and therefore, we may consider all the variables and Planck constant in Eq. (1) as dimensionless quantities. In Ref. [18], it is argued that the imaginary unit is raised to the same power as the time coordinate by performing a Wick rotation. More details about this issue are discussed in [19]. Solutions for FTSE has been investigated in many settings, including the fractional dynamics of free particles [18], and particles under the influence of $\delta$ potentials [20]. An interesting noteworthy is a mathematical correspondence between the FTSE and the fractional-time diffusion equation [18, 19, 21, 20], viewed as describing a non-Markovian process. This correspondence can be verified by replacing the real-time for the imaginary-time in the fractional diffusion equation [18]. Also, a connection between classical geometric diffusion and quantum dynamics is elucidated in Ref. [22], wherein continuous-time quantum walks are represented as quantum analogs of turbulent diffusion within comb geometry.
Recently, quantum information science has witnessed remarkable growth and development, propelled by advances in both theoretical and experimental capabilities. This promising field has revolutionized our understanding of information processing at the quantum level, such as quantum cryptography [23], quantum teleportation [24], quantum metrology [25], and quantum control [26], or quantum computing [27]. While fractional calculus offers a rich mathematical framework for describing complex phenomena, its application to quantum information problems remains relatively unexplored. In this sense, Zu and coworkers [28, 29] analyze the memory effect role of FTSE in the time-evolution of a single quantum state and quantum entanglement by considering the Jaynes-Cummings (JC) model [30]. The JC model describes the interaction between a single atom and a single light wave trapped in a cavity. This interaction exhibits fascinating quantum phenomena like Rabi oscillations and entanglement, providing insights into fundamental light-matter interactions and laying the groundwork for advancements in quantum technologies. Within the fractional scenario, the two-level system interacting with the light field is investigated in Refs. [31, 32].
In fact, the FTSE generates many undesired results, such as the non-existence of stationary energy levels, non-unitarity of the evolution, and consequently, the non-conservation of probability, as discussed in Ref. [33]. It becomes evident when we transform the FTSE in a usual Schrödinger-like equation with an effective time-dependent non-Hermitian Hamiltonian operator. It can be explicitly verified by applying the Riemann-Liouville derivative operator ${}^{\text{RL}}_{~{}~{}0}\mathcal{D}_{t}^{1-\alpha}$ on both sides of the Eq. (1), and evoking the following property of the fractional differentiation for $\alpha\in(0,1]$ [34]
${}^{\text{RL}}_{~{}~{}0}\mathcal{D}_{t}^{1-\alpha}\;{}^{\text{C}}_{0}\mathcal{D}_{t}^{\alpha}|\Psi_{\alpha}(t)\rangle=\partial_{t}|\Psi_{\alpha}(t)\rangle,$ (3)
with
[…]
Eq. (1) becomes
$i\hbar\partial_{t}|\Psi_{\alpha}(t)\rangle=\frac{i^{1-\alpha}\hbar}{\hbar_{\alpha}}\hat{H}_{\alpha}{\,}^{\text{RL}}_{~{}~{}0}\mathcal{D}_{t}^{1-\alpha}|\Psi_{\alpha}(t)\rangle.$ (5)
As mentioned before, the effective Hamiltonian is non-Hermitian, which implies a non-unitary time-evolution of the quantum state. In this sense, different proposals to map the non-unitary fractional evolution operator into a unitary one have been made [35, 33, 36, 37]. Particularly, in Ref. [37], a unitary evolution for a traceless non-Hermitian two-level system evolving under FTSE was established by applying the time-dependent non-Hermitian quantum formalism [38, 39, 40]. Essentially, the map of the non-unitary fractional time-evolution operator into a unitary one by employing non-Hermitian quantum mechanics procedures for time-dependent metrics allows for a proper quantum mechanical interpretation of the fractional dynamics. For more details about time-dependent non-Hermitian systems see Refs. [38, 41, 42, 39, 43, 44, 45, 46].
In this work, we propose a unitary fractional-time evolution of the Jaynes-Cummings (JC) model, guided by the formalism developed in [37]. In Section II, we introduce the JC model and the fractional-time evolution operator by solving the fractional time Schrödinger equation (FTSE). Additionally, we establish a connection with time-dependent non-Hermitian quantum formalism by constructing a time-dependent Dyson map, which relates to the dynamical Hilbert space metric. This approach enables us to derive an equivalent unitary time-evolution operator that describes the system’s dynamics within the conventional quantum mechanical framework. Subsequently, in Sec. III, we analyze the population inversion dynamics of an atom interacting with a coherent field. We observe collapse and revival phenomena for various values of the fractional-order parameter, $\alpha$. Following this, in Section IV, we quantify the entanglement dynamics using the von Neumann entropy, verifying that the system evolves from a separable state to an entangled one. Our conclusions follow in Sec. V.
II Fractional-time dynamics of the Jaynes-Cummings model
II.1 The model
This model serves as a paradigm for comprehending the fundamental processes governing light-matter coupling at the quantum level. The JC model describes the interaction of a two-level atom with a single quantized mode of the radiation field. In an ideal cavity QED experiment, the atom can be viewed as a two-level system ($|g\rangle$ and $|e\rangle$) coupled to a single mode of the field, and the system evolution is determined by the famous JC Hamiltonian of quantum optics. Let us consider the resonant case, where the atomic energy gap between the two-level is equal to the energy radiation field. In the interaction picture, the Hamiltonian of the JC model with the rotating-wave approximation (RWA) is described by
$\hat{H}_{\alpha}=\hbar_{\alpha}\mu_{\alpha}(\hat{\sigma}_{+}\hat{a}+\hat{\sigma}_{-}\hat{a}^{\dagger}),$ (6)
where the field is characterized by the annihilation $\hat{a}$ and creation $\hat{a}^{\dagger}$ bosonic operators satisfying the Weyl-Heisenberg algebra $[\hat{a},\hat{a}^{\dagger}]=1$. The operators $\hat{\sigma}_{+}=|e\rangle\langle g|$ and $\hat{\sigma}_{-}=|g\rangle\langle e|$ are the so-called atomic transition operators, which together with the inversion operator $\hat{\sigma}_{z}=|e\rangle\langle e|-|g\rangle\langle g|$ satisfy the $\mathfrak{su}(2)$ Lie algebra $[\hat{\sigma}_{+},\hat{\sigma}_{-}]=\hat{\sigma}_{z}$ and $[\hat{\sigma}_{z},\hat{\sigma}_{\pm}]=\pm 2\hat{\sigma}_{\pm}$. The constant $\mu_{\alpha}$ denotes the atom-field coupling coefficient which represents the strength of the atom-field coupling.
The JC model is specified via the states of both atom and field, where the basis states of the field are the number states $|n\rangle$, with $n=0,1,2,\cdots$. In this case, the bare states $|g,n\rangle$ and $|e,n\rangle$ provide a natural basis for the infinite-dimensional Hilbert space representing the atom-field interaction. The ground state corresponds to the state with the atom in the ground state $|g\rangle$ and no photons in the cavity $|0\rangle$. In this case, we have the relation
$\hat{H}_{\alpha}|g,0\rangle=0,$
which means that spontaneous absorption from the vacuum is forbidden. Furthermore, for each photon number $n$ the bare states pairs, the Hamiltonian couples the states $|e,n\rangle$ and $|g,n+1\rangle$, since
$\hat{H}_{\alpha}|e,n\rangle$ $=\hbar_{\alpha}\mu_{\alpha}\sqrt{n+1}|g,n+1\rangle$
$\hat{H}_{\alpha}|g,n+1\rangle$ $=\hbar_{\alpha}\mu_{\alpha}\sqrt{n+1}|e,n\rangle.$
Thus, the infinite-dimensional Hilbert space $\mathcal{H}$ consists of the one-dimensional subspace spanned by the ground state vector $\mathcal{H}_{\text{ground}}=\{|g,0\rangle\}$ and the mutually decoupled two-dimensional subspace $\mathcal{H}_{n}=\{|e,n\rangle,|g,n+1\rangle\}$. In other words, the Hilbert space $\mathcal{H}=L^{2}(\mathbb{R})\otimes\mathbb{C}^{2}$ decays into a direct sum of dynamically invariant subspaces
$\mathcal{H}=\mathcal{H}_{\text{ground}}\oplus\mathcal{H}_{0}\oplus\mathcal{H}_{1}\oplus\mathcal{H}_{2}\oplus\cdots\,.$
Therefore, the Hamiltonian can be decomposed as a block-diagonal matrix
$\hat{H}_{\alpha}=\hbar_{\alpha}\left[\begin{matrix}0&0_{1\times 2}&0_{1\times 2}&\cdots\\ 0_{2\times 1}&\hat{H}_{\alpha}^{(0)}&0_{2\times 2}&\cdots\\ 0_{2\times 1}&0_{2\times 2}&\hat{H}_{\alpha}^{(1)}&\cdots\\ \vdots&\vdots&\vdots&\ddots\\ \end{matrix}\right],$ (7)
where $\hat{H}_{\alpha}^{(n)}$ is the traceless $2\times 2$ matrix
$\hat{H}_{\alpha}^{(n)}=\hbar_{\alpha}\mu^{(n)}_{\alpha}\left[\begin{matrix}0&1\\ 1&0\end{matrix}\right],$ (8)
with $\mu^{(n)}_{\alpha}=\sqrt{n+1}\mu_{\alpha}$, in which represents the Hamiltonian of the system in the two-dimensional subspace $\mathcal{H}_{n}$. Next, we discuss how the system evolves in time under the FTSE.
II.2 Fractional time-evoluiton
In the fractional-time scenario, the dynamic of the system is claimed to be described by the FTSE given in Eq. (1). The formal solution of this equation can be read as $|\Psi_{\alpha}(t)\rangle=\hat{U}_{\alpha}(t)|\Psi_{\alpha}(0)\rangle$, where the system evolves from an initial state $|\Psi^{\alpha}(0)\rangle$ to the state $|\Psi^{\alpha}(t)\rangle$ through the following time-evolution operator $\hat{U}_{\alpha}(t)$,
$\hat{U}_{\alpha}(t)=E_{\alpha}\left(i^{-\alpha}\hat{H}_{\alpha}\,t^{\alpha}/\hbar_{\alpha}\right),$ (9)
that is a non-unitary operator satisfying the initial condition $\hat{U}_{\alpha}(0)=\hat{1}$. In the above equation, the function $E_{\alpha}(x)=\sum_{k=0}^{\infty}x^{k}/\Gamma(\alpha k+1)$ is identified to be the well-known one-parameter Mittag-Leffler function [1].
Once the Hamiltonian is represented in a block-diagonal form, as seen in Eq. (7), the nonunitary time-evolution operator can be represented as
$\hat{U}_{\alpha}(t)=\left[\begin{matrix}1&0_{1\times 2}&0_{1\times 2}&\cdots\\ 0_{2\times 1}&\hat{U}_{\alpha}^{(0)}(t)&0_{2\times 2}&\cdots\\ 0_{2\times 1}&0_{2\times 2}&\hat{U}_{\alpha}^{(1)}(t)&\cdots\\ \vdots&\vdots&\vdots&\ddots\\ \end{matrix}\right],$ (10)
where $\hat{U}_{\alpha}^{(n)}(t)$ is given by
$\hat{U}_{\alpha}^{(n)}(t)=\left[\begin{matrix}\mathcal{C}^{(n)}_{\alpha}(t)&i^{-\alpha}\mathcal{S}^{(n)}_{\alpha}(t)\\ i^{-\alpha}\mathcal{S}^{(n)}_{\alpha}(t)&\mathcal{C}^{(n)}_{\alpha}(t)\end{matrix}\right],$ (11)
satisfying the initial condition $\hat{U}_{\alpha}^{(n)}(0)=\hat{1}_{2\times 2}$. Here, the complex functions $\mathcal{C}^{(n)}_{\alpha}(t)$ and $\mathcal{S}^{(n)}_{\alpha}(t)$ are given in the form
$\mathcal{C}^{(n)}_{\alpha}(t)={}$ $\frac{E_{\alpha}(i^{-\alpha}\mu^{(n)}_{\alpha}t^{\alpha})+E_{\alpha}(-i^{-\alpha}\mu^{(n)}_{\alpha}t^{\alpha})}{2},$ (12a)
$\quad\mathcal{S}^{(n)}_{\alpha}(t)={}$ $\frac{E_{\alpha}(i^{-\alpha}\mu^{(n)}_{\alpha}t^{\alpha})-E_{\alpha}(-i^{-\alpha}\mu^{(n)}_{\alpha}t^{\alpha})}{2i^{-\alpha}}.$ (12b)
The nonunitary nature of time evolution when considering the FTSE might be applied to mimic the effects of the environment on quantum systems. However, this approach would be inherently heuristic, as nonunitarity leads to the non-conservation of probability. In order to establish the conventional interpretation of quantum mechanics, our aim is to map the nonunitary fractional time-evolution operator to a unitary one by employing non-Hermitian quantum mechanics techniques with time-dependent metrics. This procedure enables a proper quantum-mechanical interpretation of the fractional time description, utilizing a modified inner product as done in the non-Hermitian framework.
II.3 Unitary fractional-time evolution
Hereafter we apply the results developed in Ref. [37] for the JC model. Remarkably, within the framework of time-dependent non-Hermitian formalism [38, 40], a state undergoing nonunitary evolution, denoted as $|\Psi_{\alpha}(t)\rangle$, can be linked to a state evolving unitarily, $|\psi_{\alpha}(t)\rangle$, via the time-dependent Dyson map, expressed by the following relation:
$|\psi_{\alpha}(t)\rangle=\hat{\eta}_{\alpha}(t)|\Psi_{\alpha}(t)\rangle,$ (13)
such a map is assumed to be invertible. In what follows, from the fact that $|\psi_{\alpha}(t)\rangle=\hat{u}_{\alpha}(t)|\psi_{\alpha}(0)\rangle$, the Eq. (13) allows us to obtain the unitary time-evolution operator $\hat{u}_{\alpha}(t)$ in terms of the Dyson map $\hat{\eta}_{\alpha}(t)$ and the non-unitary time-evolution operator $\hat{U}_{\alpha}(t)$ from the equality
$\hat{u}_{\alpha}(t)=\hat{\eta}_{\alpha}(t)\hat{U}_{\alpha}(t)\hat{\eta}_{\alpha}^{-1}(0).$ (14)
Since the nonunitary time-evolution operator $\hat{U}_{\alpha}(t)$ is known, we have to specify the time-dependent Dyson map parameters for mapping the fractional dynamics in a unitary one. It shows that is possible to define a dynamical Hilbert space with a modified inner-product, defined as $\langle\Psi_{\alpha}(t)|\Psi_{\alpha}(t)\rangle_{\Theta_{\alpha}(t)}=\langle\Psi_{\alpha}(t)|\hat{\Theta}_{\alpha}(t)|\Psi_{\alpha}(t)\rangle$, where the fractional-time evolution can be seen as a unitary in according to modified inner product
$\langle\Psi_{\alpha}(t)|\Psi_{\alpha}(t)\rangle_{\Theta_{\alpha}(t)}$ $=\langle\Psi_{\alpha}(0)|\Psi_{\alpha}(0)\rangle_{\Theta_{\alpha}(0)}$
$=\langle\psi_{\alpha}(0)|\psi_{\alpha}(0)\rangle$
$=\langle\psi_{\alpha}(t)|\psi_{\alpha}(t)\rangle,$
in which $\hat{\Theta}_{\alpha}(t)=\hat{\eta}_{\alpha}^{\dagger}(t)\hat{\eta}_{\alpha}^{\,}(t)$ is the metric operator. This relation reflects that the probability conservation in the fractional-time scenario can be achieved by defining a suitable time-dependent metric with respect to which the state evolves unitarily. Also, this relation means that it is equivalent to mapping the state that evolves nonunitarily in another system that evolves unitarily with respect to the usual metric.
Indeed, the choice of time-dependent Dyson map $\hat{\eta}_{\alpha}(t)$ is not unique, and for this reason, we propose an Hermitian block-diagonal form given by
$\hat{\eta}_{\alpha}(t)=\left[\begin{matrix}1&0_{1\times 2}&0_{1\times 2}&\cdots\\ 0_{2\times 1}&\hat{\eta}_{\alpha}^{(0)}(t)&0_{2\times 2}&\cdots\\ 0_{2\times 1}&0_{2\times 2}&\hat{\eta}_{\alpha}^{(1)}(t)&\cdots\\ \vdots&\vdots&\vdots&\ddots\\ \end{matrix}\right],$ (15)
where the Dyson map acting on the two-dimensional subspace $\mathcal{H}_{n}$ is chosen as being
$\hat{\eta}_{\alpha}^{(n)}(t)=e^{\kappa_{\alpha}^{(n)}(t)}e^{\lambda_{\alpha}^{(n)}(t)\hat{\sigma}_{+}}e^{\ln{\Lambda_{\alpha}^{(n)}(t)}\hat{\sigma}_{z}/2}e^{[\lambda_{\alpha}^{(n)}(t)]^{\ast}\hat{\sigma}_{-}},$
where we assume that $\lambda_{\alpha}^{(n)}(t)\in\mathbb{C}$ and $\kappa_{\alpha}^{(n)}(t),\Lambda_{\alpha}^{(n)}(t)\in\mathbb{R}$ under the additional condition $\Lambda_{\alpha}^{(n)}(t)>0$. Moreover, we can also represent the $n$-th subspace of Dyson map as a matrix in the basis $\{|e,n\rangle,|g,n+1\rangle\}$, it yields
$\hat{\eta}_{\alpha}^{(n)}(t)=\frac{e^{\kappa_{\alpha}^{(n)}}}{\sqrt{\Lambda_{\alpha}^{(n)}}}\left[\begin{matrix}\chi_{\alpha}^{n}&\lambda_{\alpha}^{(n)}\\ [\lambda_{\alpha}^{(n)}]^{\ast}&1\end{matrix}\right],$ (16)
where we omit the time dependencies in the Dyson map parameters to clean up the notation and define the function $\chi_{\alpha}^{(n)}(t)=\Lambda_{\alpha}^{(n)}+|\lambda_{\alpha}^{(n)}|^{2}$.
In applying the results of Eqs. (11) and (16) into Eq. (14), we have the operator $\hat{u}_{\alpha}(t)$ to be represented as
$\hat{u}_{\alpha}(t)=\left[\begin{matrix}1&0_{1\times 2}&0_{1\times 2}&\cdots\\ 0_{2\times 1}&\hat{u}_{\alpha}^{(0)}(t)&0_{2\times 2}&\cdots\\ 0_{2\times 1}&0_{2\times 2}&\hat{u}_{\alpha}^{(1)}(t)&\cdots\\ \vdots&\vdots&\vdots&\ddots\\ \end{matrix}\right],$ (17)
with $\hat{u}_{\alpha}^{(n)}(t)=\hat{\eta}_{\alpha}^{(n)}(t)\hat{U}_{\alpha}^{(n)}(t)[\hat{\eta}_{\alpha}^{(n)}(0)]^{-1}$. Admitting $\hat{u}_{\alpha}(t)$ as a unitary operator, implies on $\hat{u}_{\alpha}^{(n)}(t)$ must necessarily belong to the Lie group $U(2)$, and it can be represented in the matrix form
$\hat{u}^{(n)}_{\alpha}(t)=e^{i\delta_{\alpha}^{(n)}}\left[\begin{matrix}\varpi_{\alpha,+}^{(n)}&\varpi_{\alpha,-}^{(n)}\\ -[\varpi_{\alpha,-}^{(n)}]^{\ast}&[\varpi_{\alpha,+}^{(n)}]^{\ast}\end{matrix}\right],$ (18)
with
$\varpi_{\alpha,\pm}^{(n)}(t)={}$$\pm e^{i\delta_{\alpha}^{(n)}}[\nu_{\alpha,\mp}^{(n)}]^{\ast},$ (19a)
$\delta_{\alpha}^{(n)}(t)={}$ $\frac{1}{2}\mathrm{Im}[\ln{D_{\alpha}^{(n)}}],$ (19b)
in which $D_{\alpha}^{(n)}(t)=[\mathcal{C}^{(n)}_{\alpha}]^{2}-(-1)^{-\alpha}[\mathcal{S}^{(n)}_{\alpha}]^{2}$, and the coefficients satisfying the relation $|\varpi_{\alpha,+}^{(n)}|^{2}+|\varpi_{\alpha,-}^{(n)}|^{2}=1.$ In Eq. (19), the functions $\nu_{\alpha,\pm}^{(n)}$ are given by
$\nu_{\alpha,\pm}^{(n)}(t)={}$$\pm\frac{e^{\kappa_{\alpha}^{(n)}-\kappa_{\alpha}^{(n)}(0)}}{\sqrt{\Lambda_{\alpha}^{(n)}\Lambda_{\alpha}^{(n)}(0)}}\left[\zeta_{\alpha,\pm}^{(n)}+[\lambda_{\alpha}^{(n)}]^{\ast}\xi_{\alpha,\pm}^{(n)}\right],$ (20)
with the functions $\zeta_{\alpha,\pm}^{(n)}$ e $\xi_{\alpha,\pm}^{(n)}$ defined as
$\zeta_{\alpha,+}^{(n)}(t)$ $=i^{-\alpha}\mathcal{S}_{\alpha}^{(n)}-[\lambda_{\alpha}^{(n)}(0)]^{\ast}\mathcal{C}_{\alpha}^{(n)},$ (21a)
$\zeta_{\alpha,-}^{(n)}(t)$ $=i^{-\alpha}\lambda_{\alpha}^{(n)}(0)\mathcal{S}_{\alpha}^{(n)}-\chi_{\alpha}^{(n)}(0)\mathcal{C}_{\alpha}^{(n)},$ (21b)
$\xi_{\alpha,+}^{(n)}(t)$ $=\mathcal{C}_{\alpha}^{(n)}-i^{-\alpha}[\lambda_{\alpha}^{(n)}(0)]^{\ast}\mathcal{S}_{\alpha}^{(n)},$ (21c)
$\xi_{\alpha,-}^{(n)}(t)$ $=\lambda_{\alpha}^{(n)}(0)\mathcal{C}_{\alpha}^{(n)}-i^{-\alpha}\chi_{\alpha}^{(n)}(0)\mathcal{S}_{\alpha}^{(n)}.$ (21d)
From the imposition of unitarity of $\hat{u}_{\alpha}^{(n)}$, the Dyson map parameters have the form
$\kappa_{\alpha}^{(n)}(t)={}$ $\kappa_{\alpha}^{(n)}(0)-\frac{1}{2}\mathrm{Re}[\ln{D_{\alpha}^{(n)}}],$ (22a)
$\chi_{\alpha}^{(n)}(t)={}$ $\frac{|\zeta_{\alpha,+}^{(n)}|^{2}+|\zeta_{\alpha,-}^{(n)}|^{2}+\Lambda_{\alpha}^{(n)}(0)e^{\mathrm{Re}[\ln{D_{\alpha}^{(n)}}]}}{|\xi_{\alpha,+}^{(n)}|^{2}+|\xi_{\alpha,-}^{(n)}|^{2}+\Lambda_{\alpha}^{(n)}(0)e^{\mathrm{Re}[\ln{D_{\alpha}^{(n)}}]}},$ (22b)
$\lambda_{\alpha}^{(n)}(t)={}$ $\frac{-\left[\xi_{\alpha,+}^{(n)}[\zeta_{\alpha,+}^{(n)}]^{\ast}+\xi_{\alpha,-}^{(n)}[\zeta_{\alpha,-}^{(n)}]^{\ast}\right]}{|\xi_{\alpha,+}^{(n)}|^{2}+|\xi_{\alpha,-}^{(n)}|^{2}+\Lambda_{\alpha}^{(n)}(0)e^{\mathrm{Re}[\ln{D_{\alpha}^{(n)}}]}},$ (22c)
where the time-dependent Dyson map parameter $\Lambda_{\alpha}^{(n)}=\chi_{\alpha}^{(n)}-|\lambda_{\alpha}^{(n)}|^{2}$, in which we must have $\chi_{\alpha}^{(n)}>|\lambda_{\alpha}^{(n)}|^{2}$, once we are assuming $\Lambda_{\alpha}^{(n)}$ as a positive function. Notice that these functions depend only on the fractional-time evolution parameters and on the initial values of Dyson map parameters into the functions $\xi_{\alpha,\pm}^{(n)}(t)$ and $\zeta_{\alpha,\pm}^{(n)}(t)$. For more details about those calculations see Ref. [37].
III Atomic Population Inversion: collapse and revival
The phenomenon of collapse and revival of atomic oscillations is a distinctive feature observed in the interaction of a two-level atom with a quantized electromagnetic field inside a cavity described by the JC model. When the cavity field is prepared in a coherent state and interacts with the two-level atom, the system undergoes a fascinating dynamical behavior characterized by the following stages: i) initial Rabi Oscillations: at the outset, the atom exchanges energy with the quantized field, resulting in oscillations of the atomic population between the ground and excited states. These are known as Rabi oscillations; ii) collapse: due to the quantized nature of the field and the distribution of photon number states in the coherent state, these oscillations begin to dephase. This dephasing causes the observable oscillations to diminish in amplitude, leading to a phenomenon known as the "collapse" of the Rabi oscillations; iii) revival: after a certain period, the oscillations rephase, and the atomic population oscillations reappear. This re-emergence of oscillations is termed the "revival" of the Rabi oscillations. The timing of these revivals is determined by the properties of the coherent state and the parameters of the system.
We aim to analyze this phenomenon depending on the $\alpha$ parameter in the unitary approach. In what follows, we consider the situation in which the atom is initially prepared in the excited state and the field is a coherent state, such that $|\psi_{\alpha}(0)\rangle=|e\rangle\otimes|\beta\rangle$, where $|\beta\rangle=\sum_{n=0}^{\infty}c_{n}|n\rangle$ being $c_{n}=e^{-|\beta|^{2}/2}\beta^{n}/\sqrt{n!}$ with $\beta\in\mathbb{C}$. The system evolves in according to $|\psi_{\alpha}(t)\rangle=\hat{u}_{\alpha}(t)|\psi_{\alpha}(0)\rangle$ that results in
$|\psi_{\alpha}(t)\rangle=\sum_{n=0}^{\infty}\Big{[}A_{e,n}^{\alpha}(t)|e,n\rangle+A_{g,n}^{\alpha}(t)|g,n+1\rangle\Big{]},$ (23)
where the time-dependent probability amplitudes are
$A_{e,n}^{\alpha}(t)$ $=c_{n}e^{i\delta_{\alpha}^{(n)}(t)}\varpi_{\alpha,+}^{(n)}(t),$ (24a)
$A_{g,n}^{\alpha}(t)$ $=-c_{n}e^{i\delta_{\alpha}^{(n)}(t)}[\varpi_{\alpha,-}^{(n)}(t)]^{\ast}.$ (24b)
The probability of finding the atom in the excited state with the field having $n$ photons is $P_{e,n}^{\alpha}(t)=|A_{e,n}^{\alpha}(t)|^{2}$. In contrast, the probability of finding the atom in the ground state with the field having $n+1$ photons is $P_{g,n+1}^{\alpha}(t)=|A_{g,n}^{\alpha}(t)|^{2}$. Furthermore, we can marginalize the probability over the field states, by summing over all possible photon numbers to obtain the probability of finding the atom in an excited or grounded state, which are given respectively by $P_{e}^{\alpha}(t)=\sum_{n=0}^{\infty}|A_{e,n}^{\alpha}(t)|^{2}$ and $P_{g}^{\alpha}(t)=\sum_{n=0}^{\infty}|A_{g,n}^{\alpha}(t)|^{2}$ We then calculate the population inversion of the atom along the time, which is given by the mean value of the inversion operator,
$W_{\alpha}(t)$ $=P_{e}^{\alpha}(t)-P_{g}^{\alpha}(t)$
$=\sum_{n=0}^{\infty}\left[|A_{e,n}^{\alpha}(t)|^{2}-|A_{g,n}^{\alpha}(t)|^{2}\right].$ (25)
Fig. 1 illustrates the results for the coherent state parameter $\beta=2$ and various values of the parameter $\alpha$, with initial conditions for the Dyson map set as $\kappa_{\alpha}^{(n)}(0)=0$, $\Lambda_{\alpha}^{(n)}(0)=1$, and $\lambda_{\alpha}^{(n)}(0)=0$. These conditions correspond to the scenario where $\hat{\eta}_{\alpha}(0)$ is the identity operator, ensuring that the initial states in both the Hermitian and non-Hermitian representations are identical, i.e., $|\psi_{\alpha}(0)\rangle=|\Psi_{\alpha}(0)\rangle$. For $\alpha=1.00$ (black solid line), we observe the well-known result of population inversion dynamics, as the fractional derivative reduces to the standard first-order derivative in Eq. (1). When $\alpha=0.75$ (blue dashed line), the population inversion exhibits a slightly modified behavior compared to the usual case ($\alpha=1.00$), though with fewer oscillations. In contrast, for $\alpha=0.50$ (orange dotted line) the rapid oscillations in population inversion vanish, giving rise to a more periodic behavior. Additionally, the population eventually returns to the excited state, indicating that the combination of the fractional-derivative parameter and the Dyson map introduces additional driving terms in the atom-field interaction within the unitary description.
Figure 1: Time-evolution of the atomic population inversion of the atom for different values of the fractional-order parameter $\alpha$. We plot the usual case corresponding to $\alpha=1.0$ (black solid line), $\alpha=0.75$ (blue dashed line) and $\alpha=0.5$ (orange dotted line). We start with the atom in the excited state and the field in the coherent state with $\beta=2$, and the initial values of Dyson map parameters being $\kappa_{\alpha}(0)=0$, $\Lambda_{\alpha}(0)=1$ and $\lambda_{\alpha}(0)=0$.
IV Entanglement between atom-field
The entanglement phenomenon has been discussed since the beginning of quantum mechanics [47]. It describes a scenario in which the quantum states of two or more particles become so correlated that the state of one particle cannot be described independently of the others, even when they are spatially separated. This is traditionally viewed as a manifestation of non-separability [48]. In the context of Bell nonlocality, not all entangled states violate Bell’s inequality, but any state that does violate it must be entangled. Thus, entanglement is a necessary condition for violating Bell’s inequality [49]. This violation reflects deviations in the statistical correlations of quantum states from classical expectations based on local realism [50, 51]. These concepts underscore the nonclassical nature of quantum mechanics and carry profound implications for our understanding of reality.
In this section, we explore the influence of the FTSE on quantum entanglement. Specifically, in Refs. [28, 29], the FTSE is employed to investigate entanglement within the JC model. Here, we analyze the entanglement dynamics based on the unitary dynamics associated with the FTSE, using the Dyson map as previously discussed. To quantify entanglement in pure bipartite states, the von Neumann entropy [52] can be used as a measure. It is defined as:
[…]
$\hat{\rho}(t)=\hat{u}_{\alpha}(t)|\psi^{\alpha}(0)\rangle\langle\psi^{\alpha}(0)|\hat{u}_{\alpha}^{\dagger}(t),$ (27)
we can obtain the reduced density matrix $\hat{\rho}_{\text{a}}(t)$ of the atom by tracing over the degrees of freedom of the field, which gives us
$\hat{\rho}_{\text{a}}(t)=\left[\begin{matrix}\rho_{ee}^{\alpha}(t)&\rho_{eg}^{\alpha}(t)\\ [\rho_{eg}^{\alpha}(t)]^{\ast}&\rho_{gg}^{\alpha}(t)\end{matrix}\right],$ (28)
where we use the same initial state of the previous section to obtain the matrix entries $\rho_{ee}^{\alpha}(t)=\sum_{n=0}^{\infty}|A_{e,n}^{\alpha}|^{2}$, $\rho_{gg}^{\alpha}(t)=\sum_{n=0}^{\infty}|A_{g,n}^{\alpha}|^{2}$, and $\rho_{eg}^{\alpha}(t)=\sum_{n=0}^{\infty}A_{e,n+1}^{\alpha}[A_{g,n}^{\alpha}]^{\ast}$. Taking the trace of the reduced density matrix, we verify that $\rho_{ee}^{\alpha}(t)+\rho_{gg}^{\alpha}(t)=1$ having the trace-preserving property as expected. In Fig. 2, we plot the von Neumann entropy evaluated through the Eq. (26) for different values of fractional parameter $\alpha$. We set the coherent parameter $\beta=2$ and the initial Dyson map parameters to be $\kappa_{\alpha}(0)=0$, $\Lambda_{\alpha}(0)=1$ and $\lambda_{\alpha}(0)=0$. For $\alpha=1.00$, in the well-known scenario where the field and atom are initially prepared in a separable pure state, the quantum dynamics for $t>0$ lead to an increase in marginal entropies, resulting in strong entanglement between the field and the atom. A similar behavior is observed for $\alpha=0.75$. However, for $\alpha=0.25$, the entanglement dynamics undergo a significant change. The entropy initially increases but, after some oscillations, it suddenly drops to nearly zero, indicating that the composite system returns to an almost separable state. This phenomenon suggests the occurrence of the sudden birth and death of entanglement, characterized by a rapid increase in entropy followed by a swift decrease to nearly zero.
Figure 2: Time-evolution of the von Neumann entropy for different values of the fractional-order parameter $\alpha$. We plot the usual case corresponding to $\alpha=1.0$ (black solid line), $\alpha=0.75$ (blue dashed line) and $\alpha=0.5$ (orange dotted line). We start with the atom in the excited state and the field in the coherent state with $\beta=2$, and the initial values of Dyson map parameters being $\kappa_{\alpha}(0)=0$, $\Lambda_{\alpha}(0)=1$ and $\lambda_{\alpha}(0)=0$.
V Conclusions
To summarize, we investigate the dynamics of the JC model within the fractional-time scenario, utilizing the Caputo derivative as outlined in Ref. [37]. This methodology leverages the time-dependent non-Hermitian Hamiltonian theory [38, 40], where a time-dependent Dyson map is constructed to link a dynamic Hilbert space to the time evolution governed by a time-dependent non-Hermitian Hamiltonian, which is unitary. Similarly, this approach is applied in the fractional-time scenario, as demonstrated in Ref. [37] for a traceless two-level system with a general non-Hermitian Hamiltonian. This framework addresses the issue of non-unitarity and aligns with quantum mechanical principles. To apply this unitary description to the JC model, represented in an infinite-dimensional space, we decompose it into invariant two-dimensional subspaces. We then examine how the fractional-order parameter $\alpha$ in the unitary dynamics affects the collapse and revival phenomena, which are intrinsic to the quantum nature of the electromagnetic field and absent in classical systems. Additionally, we explore the atom-field entanglement across different fractional-order parameters, revealing the potential for both the emergence and disappearance of entanglement. Entanglement, a quintessential quantum phenomenon with no classical analog, provides valuable insights into quantum mechanics and is of significant interest in quantum optics and quantum information science. Our results give some connections between fractional-time and non-Hermitian quantum mechanics. By providing this analysis, we hope our results may interest the community working on fractional-time and non-Hermitian systems by inspiring the search for their unitary description, which is generally ignored in those approaches.
Acknowledgements
The author thanks Maike A. F. dos Santos for his valuable comments and clarifications that significantly improved the work. Thanks are also due to Bárbara Amaral and the members of the Quantum Foundation Group (IFUSP) for their insightful discussions. Additionally, the author would like to acknowledge financial support from Instituto Serrapilheira, and the Pró-Reitoria de Pesquisa e Inovação (PRPI) from the Universidade de São Paulo (USP) by financial support through the Programa de Estímulo à Supervisão de Pós-Doutorandos por Jovens Pesquisadores.